Tin(II) hydroxide has a Ksp value of 1.4 x 10-28. Comparing this to my experimentally determined Ksp of 7.962 x 10-15 for the salt nickel(II) hydroxide, which salt is less soluble? Explain.
Answer #1

Step 1: find solubility of Sn(OH)2

At equilibrium:

Sn(OH)2 <----> Sn2+ + 2 OH-

   s 2s

Ksp = [Sn2+][OH-]^2

1.4*10^-28=(s)*(2s)^2

1.4*10^-28= 4(s)^3

s = 3.271*10^-10 M

Step 2: find solubility of Ni(OH)2

At equilibrium:

Ni(OH)2 <----> Ni2+ + 2 OH-

   s 2s

Ksp = [Ni2+][OH-]^2

7.962*10^-15=(s)*(2s)^2

7.962*10^-15= 4(s)^3

s = 1.258*10^-5 M

Solubility is more for Ni(OH)2.

So, Ni(OH)2 is more soluble

Answer: Sn(OH)2 is less soluble